# Sylvester domains and pro-$p$ groups

### v1

@@ 4. Mild flag pro-$p$ groups
It is directly verified that $w(a)$ is a filtration function in the sense of [Laz65, Def. I.2.1.1] (or a valuation in the sense of [EJZ13, Sec. 2.2]), that is, $w$ satisfies:
@@
In addition to the above, $w$ satisfies a stronger version of (iii) called the strong ultrametric inequality: if moreover $w(a)\neq w(b)$, then $w(a+b)=\max\{w(a),w(b)\}$. Following [Laz65, I.2.2.1], we shall say that $w$ is a valuation if $w$ satisfies a stronger version of (iv): $w(ab)=w(a)w(b)$ for any pair $a,b\in S$. This is the definition also adopted in [Lab06].
@@
From the homeomorphism $G\sim N\times G/N$, we obtain that $S$ is the free profinite $R$-module on the profinite space $G/N\simeq\{1,g,g^{2},\ldots\}$. Through a base change, we get that the set of powers of $s=g-1$ also form a topological basis for $S$ over $R$, from which we can identify it as a topological abelian group with the group of formal power series in $s$ over $R$. Hence, it suffices to show that the formula (1) defining the multiplication always converges. For this, it is also sufficient to check that $\delta$ is topologically nilpotent in the induced filtration of $R$, that is, $\delta^{k}$ converges to zero pointwise. Observe that $I_{G}$ and $R$ are $\sigma$-invariant because $\sigma$ is an inner automorphism of $S$ and $N$ is normal in $G$.
@@
To describe some cases in which $w(a)$ is actually a valuation, we will make use of the following definition given in [For11, Def. 1.1]:
@@
To highlight the choice of filtration on $\mathbb{F}_{p}[\![F]\!]$, one says in that case that $G$ is a mild pro-$p$ group with respect to the $p$-Zassenhaus filtration, that is, the filtration by dimension series mod $p$ – cf. [MPQT22, Def. 4.4], [For11, Lem. 1.3 and Rem. 1.6] and [Lab06, Def. 1.1]. We will concern ourselves solely with mild groups with respect to this filtration.
@@
(4) $[x_{i},\underbrace{g,\ldots,g}_{a_{i}\text{ times}}]=h_{i}\text{ for }h_{i}\in\Phi(\widetilde{N})\text{ and }1\leq i\leq l\,,$
@@
There is an $\mathbb{F}_{p}$-vector space decomposition of $H^{1}(G,\mathbb{F}_{p})\simeq V\oplus W$ such that $V=\langle\gamma\rangle$ and $W=\langle\eta_{i}\mid 1\leq i\leq n\rangle$, and the identities in (4) show that the restriction of the cup product to $V\otimes W\to H^{2}(G,\mathbb{F}_{p})$ is surjective and that $\gamma\cup\gamma=0$ by [NSWa08, Prop. 3.9.13]. Hence, $G$ satisfies the cup-product criterion of [MPQT22, Prop. 5.8] to being a mild group (cf. [For11, Sec. 6]). Later on we shall see that mild flag presentations are also strongly free presentations, and can be used to describe the graded ring $\operatorname{Gr}(\mathbb{F}_{p}[\![G]\!])$.
@@
Since the $G/N$-action on the finite module $\mathbb{F}_{p}[\![G/N]\!]/(d_{i})$ is unipotent, there exists some positive $a_{i}$ such that $[z_{i},\underbrace{g,\ldots,g}_{a_{i}\text{ times}}]$ is a fixed point, or equivalently, there exists $h_{i}$ in $\Phi(N)$ such that
(6) $[z_{i},\underbrace{g,\ldots,g}_{a_{i}+1\text{ times}}]=h_{i}\,.$
@@
Conversely, every pro-$p$ group $G$ with a flag presentation is also free-by-$\mathbb{Z}_{p}$. If $N$ is the normal subgroup of $G$ generated by the $x_{i}$, then $G/N\simeq\mathbb{Z}_{p}$ so that $G\simeq N\rtimes\mathbb{Z}_{p}$. Taking the non-negative powers of $g$ as a continuous section $G/N\to G$, the Reidemeister-Schreier rewriting process shows that $N$ has a presentation with a generating set converging to $1$ ([RZ10, Sec. 2.4]) given by all the elements $x_{i,k}=[x_{i},\underbrace{g,\ldots,g}_{k\text{ times}}]$ and relations
@@
We now turn to the problem of showing that mild flag presentations of $G$ are strongly free, and to characterize the graded ring of $\mathbb{F}_{p}[\![G]\!]$. We recall that $\mathfrak{L}=\bigoplus_{i\geq 1}D_{i}(F)/D_{i+1}(F)$ has the structure of a free restricted Lie algebra over $\mathbb{F}_{p}$ in the sense of [Jac79, Sec. V.7]. A free generating set of this Lie algebra is the image $\xi_{1},\ldots,\xi_{n},\gamma$ of $x_{1},\ldots,x_{n},g$, and we can identify $\operatorname{Gr}(\mathbb{F}_{p}[\![F]\!])$ with the universal restricted enveloping algebra $\mathcal{U}_{\mathfrak{L}}$ of $\mathfrak{L}$ ([Laz54, Thm. 6.5] and [Laz65, Thm. A.3.5]). If the elements $\rho_{1},\ldots,\rho_{l}$ associated to a presentation of $G$ lie in $\mathfrak{L}$ and $\mathfrak{r}$ is the restricted ideal of $\mathfrak{L}$ they generate, then for $J$ the ideal of $\operatorname{Gr}(\mathbb{F}_{p}[\![F]\!])$ generated by the $\rho_{i}$ the $\operatorname{Gr}(\mathbb{F}_{p}[\![F]\!])/J$-modules $J/J\operatorname{Gr}(I)$ and $M=\mathfrak{r}/[\mathfrak{r},\mathfrak{r}]+\mathfrak{r}^{[p]}$ are isomorphic. We recall Lazard’s elimination theorem for free (unrestricted) Lie algebras:
Theorem 4.8 ([Bou89, Prop. 10]).
@@
From this, we are able to deduce a restricted variant of [Lab06, Thm 3.3], keeping the above notation and using the fact that the universal restricted enveloping algebra of $\mathfrak{L}/\mathfrak{h}$ is a polynomial algebra generated by $\gamma$.
@@
If $\rho_{1},\ldots,\rho_{l}$ are homogeneous elements of $\mathfrak{h}$ with respect to the canonical grading of $\mathfrak{L}$ which are linearly independent over $\mathbb{F}_{p}[\gamma]$ modulo $[\mathfrak{h},\mathfrak{h}]+\mathfrak{h}^{[p]}$, then they are strongly free. Moreover, the restricted Lie algebra $\mathfrak{h}/(\rho_{1},\ldots,\rho_{l})$ is free, so that
@@
with $k\geq 0$ and $1\leq i\leq l$. Hence, it suffices to show that these elements are part of a free restricted generating set of $\mathfrak{h}$, for which it is enough to show that the $\rho_{1},\ldots,\rho_{l}$ are $\mathbb{F}_{p}[\gamma]$-independent modulo $[\mathfrak{h},\mathfrak{h}]+\mathfrak{h}^{[p]}$. Since this holds by assumption, we are done. Moreover, as the $\rho_{i,k}$ form part of a free restricted basis of $\mathfrak{h}$, the quotient $\mathfrak{h}/(\rho_{1},\ldots,\rho_{l})$ is a free restricted Lie algebra, so the latter part follows. ∎
@@
The $n$ elements $\rho_{i}=[\xi_{i},\gamma]$ are strongly free, as so is, modulo $[\mathfrak{h},\mathfrak{h}]+\mathfrak{h}^{[p]}$, any $\mathbb{F}_{p}$-linearly independent subset of the $\mathbb{F}_{p}$-subspace that they span.
@@
The first part follows from Theorem 4.9 and Proposition 4.10. For the second part, note that an $\mathbb{F}_{p}$-linear independent subset of their $\mathbb{F}_{p}$-span will remain $\mathbb{F}_{p}[\gamma]$-linear independent modulo $[\mathfrak{h},\mathfrak{h}]+\mathfrak{h}^{[p]}$. ∎
@@
The $l$ relations in (4) of a mild flag presentation lie in the $\mathbb{F}_{p}$-span of $[\xi_{i},\gamma]$ modulo $[\mathfrak{h},\mathfrak{h}]+\mathfrak{h}^{[p]}$ in the language of Corollary 4.11. This can be seen by observing that: (1) the $\rho_{i}$ have degree two and (2) writing them in the basis of $D_{2}(F)/D_{3}(F)$ given by the commutators $[x_{i},x_{j}]$, $[x_{i},g]$ and possibly the squares $x_{i}^{2}$ and $g^{2}$ if $p=2$ for $1\leq i<j\leq n$, the image of $\Phi(\widetilde{N})$ in this quotient does not include expressions involving $[x_{i},g]$ and $g^{2}$. The last part follows from the universal property of restricted enveloping algebras and the identification $\operatorname{Gr}(\mathbb{F}_{p}[\![G]\!])\simeq\operatorname{Gr}(\mathbb{F}_{p}[\![F]\!])/J=\mathcal{U}_{\mathfrak{L}/\mathfrak{r}}$ of Proposition 4.4. ∎
@@
this follows from the finitely generated case of [Laz65, Prop. II.3.1.4] by taking the inverse limit of finite sets of generators. Moreover, as a subring of $S=\mathbb{F}_{p}[\![G]\!]$, it inherits the valuation $w$. Since this ring has a simple description by power series, it would be desirable to have the property that
@@
where $\operatorname{Mon}(\{X_{i,j}\})$ denotes the set of all noncommutative monomials in the $X_{i,j}$. However, this is not true for arbitrary valuations on Magnus algebras, and the valuations on $R$ that do satisfy (7) are called weight functions with respect to the basis $\{X_{i,j}\}$ in [EJZ13, Sec. 2.4]. By [EJZ13, Prop. 3.2], $w$ is a weight function with respect to $\{X_{i,j}\}$ if and only if the restricted $\mathbb{F}_{p}$-Lie algebra generated by $\overline{X_{i,j}}$ is free in this basis. This is exactly what we’ve shown in Proposition 4.10, hence:
@@ 5. The embeddings of the group algebra
The map $\iota$ is an embedding of topological rings that preserves the valuation.
@@
Both $\sigma$ and $\delta$ extend from $R$ to $\Lambda[\![t]\!]$ in such a way that $\sigma(t)=t$, $\delta(t)=0$,
[…]
@@
$\delta(a^{-1})=-\sigma(a)^{-1}\delta(a)a^{-1}=-\sigma(a)ba^{-1}t\,.\qed$
@@
Let $J$ be the two-sided ideal of $\mathbb{D}$ generated by $t$ and $s$. By the construction of the embedding, the restriction of the $J$-adic topology to $S$ as a subspace of $\mathbb{D}$ coincides with the topology given by the powers of the augmentation ideal on $S$. Since $J$ is contained in the right ideal generated by $t$ and $s$, we have the identities:
@@
[…]
@@
To show that it is a domain, note that the product of homogeneous elements in $\mathbb{D}$ is non-zero in the graded ring, and see, for instance, [NO82, Cor. D.IV.5] for lifting the Noetherian property from $\operatorname{Gr}\mathbb{D}$ to $\mathbb{D}$. Then, the existence of $\mathcal{Q}$ follows from [Rob01, Thm. 2.1.15]. ∎
@@ 6.1. Filtered modules
Since we must state general results about filtered rings that will apply both to $S$ and $\mathbb{D}$, we will denote a general filtered ring by $\Omega$ with a filtration by $\Omega$-submodules $\Omega_{k}$, $k\in\mathbb{Z}$. Following the terminology in [JZ20, Sec. 4.1], a filtration $\{M_{k}\}$ on a $\Omega$-module $M$ is:
@@
If $\Omega$ is bounded and complete as a left $\Omega$-module and $M$ is a finitely generated and separated $\Omega$-module, then $M$ is also complete ([JZ20, Prop. 4.2]).
@@
Then, for $M=\Omega$, we have that $\Omega(V)$ is a free $\Omega$-module satisfying the following universal property: identifying $V$ with the set of elements $(0,\ldots,1,\ldots)\in\Omega(V)$, any function $f\colon V\to M$ from $V$ to a filtered $\Omega$-module $N$ such that $f(v)\in N_{v}$ extends uniquely to a strict homomorphism $\varphi\colon\Omega(V)\to N$. Moreover, observe that $\Omega(V)$ is bounded as a filtered $\Omega$-module if and only if $V$ is bounded as a multiset of integers.
@@
Proposition 6.1 ([JZ20, Prop. 4.5 and Cor. 4.6]).
Let $M$ be a complete and bounded filtered $\Omega$-module for a complete and bounded filtered ring $\Omega$. Then, there are a bounded multisets of integers $V_{i}$ and a strict exact sequence of filtered $\Omega$-modules
@@
While the modules $\widehat{\Omega(V)}$ are no longer free over $\Omega$ even if $V$ is bounded, if $\Omega$ is pseudocompact in the sense of [Bru66] then they are still flat with respect to complete tensor product
@@
of filtered $\Omega$-modules $M$ and $N$ ([Bru66, Cor. 1.3 and Lem. 2.1]), for in this case $\widehat{\Omega(V)}$ can be identified as a topological $\Omega$-module with the free pseudocompact $\Omega$-module $\Omega[\![\bm{V}]\!]$ on set $\bm{V}$ ([Bru66, p. 444]). This applies in particular to $\Omega=S$ and to $\Omega=\mathbb{D}$.
Let us denote by $\mathfrak{Tor}_{i}^{\Omega}(-,-)$ the $i$-derived bifunctor obtained from the complete tensor product, hereby referred to as the $i$-th continuous Tor group, in order to differentiate it from the usual $i$-derived functor $\operatorname{Tor}_{i}^{\Omega}(-,-)$ obtained from the usual tensor product $-\otimes_{\Omega}-$. Therefore, we already know that $\mathfrak{Tor}_{i}^{\Omega}(\widehat{\Omega(V)},M)=\mathfrak{Tor}_{i}^{\Omega}(N,\widehat{\Omega(V)})=0$ for any $i>0$ and $\Omega$-modules $M$ and $N$. The following proposition gives us a way to compare both functors:
Proposition 6.2 ([JZ20, Prop. 4.9]).
@@
This proposition implies the following two results, well known to the specialist, but for which we have found no reference in the literature. Observe that the isomorphisms of Proposition 6.2 are all functorial in $M$. First we show a partial form of [Lan02, Lem. XX.6.3]:
@@
satisfy $\operatorname{Tor}_{i}^{\Omega}(M,X_{j})=0$ for every $j$ and every $21\leq i\leq k-$ 2 with $k\geq 2$. Then,
@@
Observe that $Z_{2}$ also satisfies $\operatorname{Tor}_{i}^{\Omega}(M,Z_{2})=0$ now for every $31\leq i\leq k-$ 3, since we have the exact sequence
@@
for $i\leq k-1$, where this complex is obtained by applying $M\otimes_{\Omega}-$ to the sequence (italic-$\star$6.4). In particular, one has $\mathfrak{Tor}_{i}^{\Omega}(M,N)\simeq\operatorname{Tor}_{i}^{\Omega}(M,N)$ for $i\leq k-1$.
@@
This can be thought as a “partial” analogue of how one may compute derived functors through acyclic resolutions (cf. [Lan02, Thm. XX.6.2]). We consider another resolution of $N$ by projective $\Omega$-modules:
@@
and we show there exists a morphism of complexes $P_{i}\to\widehat{\Omega(V_{i})}$ extending the identity on $N$ and satisfying
[…]
{tikzcd}
@@
shows that $\operatorname{Tor}_{i}^{\Omega}(M,X_{j})=0$ for every $21\leq i\leq k-$ 2. Hence, by Lemma 6.3, we know that the diagram
{tikzcd}
has an exact top row and each column is a short exact sequence because of the vanishing of $\operatorname{Tor}_{1}(M,\widehat{\Omega(V_{i})})$. Then, applying the Snake Lemma to each column one finds the exact sequences
@@
We also state the following result of [JZ20], which we will use to lift vanishing results about the $\operatorname{Tor}$ groups over $\operatorname{Gr}S$ to those over $S$:
Theorem 6.5 ([JZ20, Thm. 4.10]).

### v2

@@ 4. Mild flag pro-$p$ groups
It is directly verified that $w(a)$ is a filtration function in the sense of [31, Def. I.2.1.1] (or a valuation in the sense of [12, Sec. 2.2]), that is, $w$ satisfies:
@@
The conditions above also imply a stronger version of (iii) called the strong ultrametric inequality: if moreover $w(a)\neq w(b)$, then $w(a+b)=\max\{w(a),w(b)\}$. Following [31, I.2.2.1], we shall say that $w$ is a valuation if $w$ satisfies a stronger version of (iv): $w(ab)=w(a)w(b)$ for any pair $a,b\in S$. This is the definition also adopted in [28].
@@
From the homeomorphism $G\sim N\times G/N$, we obtain that $S$ is the free profinite $R$-module on the profinite space $G/N\simeq\{1,g,g^{2},\ldots\}$. Through a base change, we get that the set of powers of s=g1s=g-1 also form a topological basis for $S$ over $R$, from which we can identify it as a topological abelian group with the group of formal power series in $s$ over $R$. Hence, it suffices to show that the formula (1) defining the multiplication always converges. For this, it is also sufficient to check that $\delta$ is topologically nilpotent in the induced filtration of $R$, that is, $\delta^{k}$ converges to zero pointwise. Observe that $I_{G}$ and $R$ are $\sigma$-invariant because $\sigma$ is an inner automorphism of $S$ and $N$ is normal in $G$.
@@
To describe some cases in which $w(a)$ is actually a valuation, we will make use of the following definition given in [13, Def. 1.1]:
@@
To highlight the choice of filtration on $\mathbb{F}_{p}[\![F]\!]$, one says in that case that $G$ is a mild pro-$p$ group with respect to the $p$-Zassenhaus filtration, that is, the filtration by dimension series mod $p$ – cf. [34, Def. 4.4], [13, Lem. 1.3 and Rem. 1.6] and [28, Def. 1.1]. We will concern ourselves solely with mild groups with respect to this filtration.
@@
(4) $x_{i},\underbrace{g,\ldots,g}_{a_{i}\text{ times}}]=h_{i}\text{ for }h_{i}\in\Phi(\widetilde{N})\text{ and }1\leq i\leq l\,$,
@@
There is an $\mathbb{F}_{p}$-vector space decomposition of $H^{1}(G,\mathbb{F}_{p})\simeq V\oplus W$ such that $V=\langle\gamma\rangle$ and $W=\langle\eta_{i}\mid 1\leq i\leq n\rangle$, and the identities in (4) show that the restriction of the cup product to $V\otimes W\to H^{2}(G,\mathbb{F}_{p})$ is surjective and that $\gamma\cup\gamma=0$ by [36, Prop. 3.9.13]. Hence, $G$ satisfies the cup-product criterion of [34, Prop. 5.8] to being a mild group (cf. [13, Sec. 6]). Later on we shall see that mild flag presentations are also strongly free presentations, and can be used to describe the graded ring $\operatorname{Gr}(\mathbb{F}_{p}[\![G]\!])$.
@@
Since the $G/N$-action on the finite module $\mathbb{F}_{p}[\![G/N]\!]/(d_{i})$ is unipotent, there exists some positive $a_{i}$ such that $z_{i},\underbrace{g,\ldots,g}_{a_{i}\text{ times}}]$ is a fixed point, or equivalently, there exists $h_{i}$ in $\Phi(N)$ such that
(6) $z_{i},\underbrace{g,\ldots,g}_{a_{i}+1\text{ times}}]=h_{i}\,.$
@@
Conversely, every pro-$p$ group $G$ with a flag presentation is also free-by-$\mathbb{Z}_{p}$. If $N$ is the normal subgroup of $G$ generated by the $x_{i}$, then $G/N\simeq\mathbb{Z}_{p}$ so that $G\simeq N\rtimes\mathbb{Z}_{p}$. Taking the non-negative powers of $g$ as a continuous section $G/N\to G$, the Reidemeister-Schreier rewriting process shows that $N$ has a presentation with a generating set converging to $1$ ([41, Sec. 2.4]) given by all the elements $x_{i,k}=[x_{i},\underbrace{g,\ldots,g}_{k\text{ times}}]$ and relations
@@
We now turn to the problem of showing that mild flag presentations of $G$ are strongly free, and to characterize the graded ring of $\mathbb{F}_{p}[\![G]\!]$. We recall that $\mathfrak{L}=\bigoplus_{i\geq 1}D_{i}(F)/D_{i+1}(F)$ has the structure of a free restricted Lie algebra over $\mathbb{F}_{p}$ in the sense of [18, Sec. V.7]. A free generating set of this Lie algebra is the image $\xi_{1},\ldots,\xi_{n},\gamma$ of $x_{1},\ldots,x_{n},g$, and we can identify $\operatorname{Gr}(\mathbb{F}_{p}[\![F]\!])$ with the universal restricted enveloping algebra $\mathcal{U}_{\mathfrak{L}}$ of $\mathfrak{L}$ ([30, Thm. 6.5] and [31, Thm. A.3.5]). If the elements $\rho_{1},\ldots,\rho_{l}$ associated to a presentation of $G$ lie in $\mathfrak{L}$ and $\mathfrak{r}$ is the restricted ideal of $\mathfrak{L}$ they generate, then for $J$ the ideal of $\operatorname{Gr}(\mathbb{F}_{p}[\![F]\!])$ generated by the $\rho_{i}$ the $\operatorname{Gr}(\mathbb{F}_{p}[\![F]\!])/J$-modules $J/J\operatorname{Gr}(I)$ and $M=\mathfrak{r}/[\mathfrak{r},\mathfrak{r}]+\mathfrak{r}^{[p]}$ are isomorphic. We recall Lazard’s elimination theorem for free (unrestricted) Lie algebras:
Theorem 4.8 ([7, Prop. 10]).
@@
From this, we are able to deduce a restricted variant of [28, Thm 3.3], keeping the above notation and using the fact that the universal restricted enveloping algebra of $\mathfrak{L}/\mathfrak{h}$ is a polynomial algebra generated by $\gamma$.
@@
If $\rho_{1},\ldots,\rho_{l}$ are homogeneous elements of $\mathfrak{h}$ with respect to the canonical grading of $\mathfrak{L}$ which are linearly independent over $\mathbb{F}_{p}[\gamma]$ modulo $\mathfrak{h},\mathfrak{h}]+\mathfrak{h}^{[p]}$, then they are strongly free. Moreover, the restricted Lie algebra $\mathfrak{h}/(\rho_{1},\ldots,\rho_{l})$ is free, so that
@@
with $k\geq 0$ and $1\leq i\leq l$. Hence, it suffices to show that these elements are part of a free restricted generating set of $\mathfrak{h}$, for which it is enough to show that the $\rho_{1},\ldots,\rho_{l}$ are $\mathbb{F}_{p}[\gamma]$-independent modulo $\mathfrak{h},\mathfrak{h}]+\mathfrak{h}^{[p]}$. Since this holds by assumption, we are done. Moreover, as the $\rho_{i,k}$ form part of a free restricted basis of $\mathfrak{h}$, the quotient $\mathfrak{h}/(\rho_{1},\ldots,\rho_{l})$ is a free restricted Lie algebra, so the latter part follows. ∎
@@
The $n$ elements $\rho_{i}=[\xi_{i},\gamma]$ are strongly free, as so is, modulo $\mathfrak{h},\mathfrak{h}]+\mathfrak{h}^{[p]}$, any $\mathbb{F}_{p}$-linearly independent subset of the $\mathbb{F}_{p}$-subspace that they span.
@@
The first part follows from Theorem 4.9 and Proposition 4.10. For the second part, note that an $\mathbb{F}_{p}$-linear independent subset of their $\mathbb{F}_{p}$-span will remain $\mathbb{F}_{p}[\gamma]$-linear independent modulo $\mathfrak{h},\mathfrak{h}]+\mathfrak{h}^{[p]}$. ∎
@@
[…]
@@
this follows from the finitely generated case of [31, Prop. II.3.1.4] by taking the inverse limit of finite sets of generators. Moreover, as a subring of $S=\mathbb{F}_{p}[\![G]\!]$, it inherits the valuation $w$. Since this ring has a simple description by power series, it would be desirable to have the property that
@@
where $\operatorname{Mon}(\{X_{i,j}\})$ denotes the set of all noncommutative monomials in the $X_{i,j}$. However, this is not true for arbitrary valuations on Magnus algebras, and the valuations on $R$ that do satisfy (7) are called weight functions with respect to the basis $\{X_{i,j}\}$ in [12, Sec. 2.4]. By [12, Prop. 3.2], $w$ is a weight function with respect to $\{X_{i,j}\}$ if and only if the restricted $\mathbb{F}_{p}$-Lie algebra generated by $\overline{X_{i,j}}$ is free in this basis. This is exactly what we’ve shown in Proposition 4.10, hence:
@@ 5. The embeddings of the group algebra
The map $\iota$ is an embedding of topological rings that preserves the valuation (i.e. an isometric embedding).
@@
Both $\sigma$ and $\delta$ extend from $R$ to $\Lambda[\![t]\!]$ in such a way that $\sigma$ is an automorphism, $\delta$ is a $\sigma$-derivation, $\sigma(t)=t$, $\delta(t)=0$,
$\sigma(a_{i})=\frac{\iota(\sigma(y_{i}))}{t^{w_{i}}}\quad\text{and}\quad\delta(a_{i})=\frac{\iota(\delta(y_{i}))}{t^{w_{i}}}\,.$
@@
$\delta(a^{-1})=-\sigma(a)^{-1}\delta(a)a^{-1}=-\sigma(a)^{-1}ba^{-1}t\,.\qed$
@@
Let $J$ be the two-sided ideal of $\mathbb{D}$ generated by $t$ and $s$. By the construction of the embedding, the restriction of the $J$-adic topology to $S$ as a subspace of $\mathbb{D}$ coincides with the topology given by the powers of the augmentation ideal on $S$. Since $\delta(\mathcal{D}[\![t]\!])\subseteq t\mathcal{D}[\![t]\!]$, the skew relation defining $s$ gives us
$\mathcal{D}[\![t]\!]s\subseteq s\mathcal{D}[\![t]\!]+t\mathcal{D}[\![t]\!]\,$,
and therefore $J$ is contained in the right ideal generated by $t$ and $s$. Inductively we have the identities:
@@
$\overline{\sigma}\colon J^{n}/J^{n+1}\to J^{n}/J^{n+1}\quad\text{and}\quad\overline{\delta}\colon J^{n}/J^{n+1}\to J^{n+1}/J^{n+2}\,$
@@
To show that it is a domain, note that the product of homogeneous elements in $\mathbb{D}$ is non-zero in the graded ring, and see, for instance, [35, Cor. D.IV.5] for lifting the Noetherian property from $\operatorname{Gr}\mathbb{D}$ to $\mathbb{D}$. Then, the existence of $\mathcal{Q}$ follows from [43, Thm. 2.1.15]. ∎
@@ 6.1. Filtered modules
Since we must state general results about filtered rings that will apply both to $S$ and $\mathbb{D}$, we will denote a general filtered ring by $\Omega$ with a filtration by $\Omega$-submodules $\Omega_{k}$, $k\in\mathbb{Z}$. Following the terminology in [24, Sec. 4.1], a filtration $\{M_{k}\}$ on a $\Omega$-module $M$ is:
@@
If $\Omega$ is bounded and complete as a left $\Omega$-module and $M$ is a finitely generated and separated $\Omega$-module, then $M$ is also complete ([24, Prop. 4.2]).
@@
Then, for $M=\Omega$, we have that $\Omega(V)$ is a free $\Omega$-module satisfying the following universal property: identifying $V$ with the set of elements $(0,\ldots,1,\ldots)\in\Omega(V)$, any function $f\colon V\to N$ from $V$ to a filtered $\Omega$-module $N$ such that $f(v)\in N_{v}$ extends uniquely to a strict homomorphism $\varphi\colon\Omega(V)\to N$. Moreover, observe that $\Omega(V)$ is bounded as a filtered $\Omega$-module if and only if $V$ is bounded as a multiset of integers.
@@
Proposition 6.1 ([24, Prop. 4.5 and Cor. 4.6]).
Let $M$ be a complete and bounded filtered $\Omega$-module for a complete and bounded filtered ring $\Omega$. Then, there are bounded multisets of integers $V_{i}$ and a strict exact sequence of filtered $\Omega$-modules
@@
While the modules $\widehat{\Omega(V)}$ are no longer free over $\Omega$ even if $V$ is bounded, if $\Omega$ is pseudocompact in the sense of [8] then they are still flat with respect to complete tensor product
@@
of filtered $\Omega$-modules $M$ and $N$ ([8, Cor. 1.3 and Lem. 2.1]), for in this case $\widehat{\Omega(V)}$ can be identified as a topological $\Omega$-module with the free pseudocompact $\Omega$-module $\Omega[\![\bm{V}]\!]$ on set $\bm{V}$ ([8, p. 444]). This applies in particular to $\Omega=S$ and to $\Omega=\mathbb{D}$.
Let us denote by $\mathfrak{Tor}_{i}^{\Omega}(-,-)$ the $i$-derived bifunctor obtained from the complete tensor product, hereby referred to as the $i$-th continuous Tor group, in order to differentiate it from the usual $i$-derived functor $\operatorname{Tor}_{i}^{\Omega}(-,-)$ obtained from the usual tensor product $\otimes_{\Omega}-$. Therefore, we already know that $\mathfrak{Tor}_{i}^{\Omega}(\widehat{\Omega(V)},M)=\mathfrak{Tor}_{i}^{\Omega}(N,\widehat{\Omega(V)})=0$ for any $i>0$ and $\Omega$-modules $M$ and $N$. The following proposition gives us a way to compare both functors:
Proposition 6.2 ([24, Prop. 4.9]).
@@
This result implies the following proposition, well known to the specialist, but for which we have found no reference in the literature. Observe that the isomorphisms of Proposition 6.2 are all functorial in $M$. First we show a partial form of [29, Lem. XX.6.3]:
@@
satisfy $\operatorname{Tor}_{i}^{\Omega}(M,X_{j})=0$ for every $j$ and every $1\leq i\leq k-2$ with $k\geq 2$. Then,
@@
Observe that $Z_{2}$ also satisfies $\operatorname{Tor}_{i}^{\Omega}(M,Z_{2})=0$ now for every $1\leq i\leq k-3$, since we have the exact sequence
@@
for $i\leq k-1$, where this complex is obtained by applying $M\otimes_{\Omega}$- to the sequence ($\star$6.4). In particular, one has $\mathfrak{Tor}_{i}^{\Omega}(M,N)\simeq\operatorname{Tor}_{i}^{\Omega}(M,N)$ for $i\leq k-1$.
@@
This can be thought as a “partial” analogue of how one may compute derived functors through acyclic resolutions (cf. [29, Thm. XX.6.2]). We consider another resolution of $N$ by projective $\Omega$-modules:
$\star\star$ $\cdots\to P_{i}\to\cdots\to P_{1}\to P_{0}\to N\to 0\,$,
and we will show that there exists a morphism of complexes $P_{i}\to\widehat{\Omega(V_{i})}$ extending the identity on $N$ and satisfying
Since $\operatorname{Tor}_{i}^{\Omega}(M,N)$ doesn’t depend on the choice of projective resolution ($\star\star$6.1), we can inductively choose the $P_{i}$ such that we have a commuting diagram with surjective columns:
@@
shows that $\operatorname{Tor}_{i}^{\Omega}(M,X_{j})=0$ for every $1\leq i\leq k-2$. Hence, by Lemma 6.3, we know that the diagram
[…]
@@
We also state the following result of [24], which we will use to lift vanishing results about the $\operatorname{Tor}$ groups over $\operatorname{Gr}S$ to those over $S$:
Theorem 6.5 ([24, Thm. 4.10]).